# The potential applications of muography to revealing sea shipwrecks

### v3

@@ Abstract
Muon imaging, a non-invasive technique leveraging naturally occurring cosmic muons, has emerged as a promising tool for exploring underwater objects, including shipwrecks. This article examines the potential of muon imaging to investigate the contents of wrecked ships in the Baltic Sea and other marine environments. By analyzing the principles of muon tomography, its applications in underwater archaeology, and its advantages over traditional methods, we highlight its potential to reveal hidden details within shipwrecks while addressing challenges posed by the marine environment. These wrecks pose significant environmental risks due to their hazardous contents, such as explosives and crude oil products, making their detection and monitoring a matter of environmental and safety concern.
Accurate modeling and imaging of such wrecks are therefore crucial for assessing potential dangers and mitigating environmental impacts. In this work, we adopted an approach similar to that used in underground muon experiments.
Cosmic muons were propagated through 35 m of seawater using GEANT4. The muons were tracked from the sea surface to a depth of 35 m, where their kinematics and energy deposition were recorded in a virtual detector. The resulting distributions were then tabulated and used as input for the PrimaryGeneratorAction module. This approach enabled the generation of a realistic underwater muon flux over a $20\times 20$ m2 surface at 35 meters depth, which was subsequently employed to simulate different shipwreck filling scenarios. Assuming an exposure time of one week for a wreck located at a depth of 50 meters, our simulations demonstrate that muon imaging can resolve density contrasts sufficiently to distinguish between water, oil, and high-density materials, underscoring its feasibility as a practical tool for underwater hazard assessment.
@@ 1 Introduction
The deep sea holds countless shipwrecks, fragile time capsules that preserve traces of past civilizations. Yet investigating their contents is difficult: direct access risks both damage to the wreck and the loss of valuable archaeological context.
A promising, non-invasive solution comes from particle physics: muon imaging. Also known as muon tomography or radiography, this technique exploits naturally occurring cosmic-ray muons, highly penetrating particles that can reveal the internal structure of large, dense objects. Initially developed for applications such as scanning pyramids and monitoring nuclear reactors, muon imaging has since expanded to geophysics, archaeology, and civil engineering [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14].
Here, we consider its application to underwater environments. By imaging shipwrecks with muons, researchers may identify and study their internal structures and contents without excavation, offering a new way to preserve cultural heritage while unlocking historical and scientific insights.
@@
Underwater shipwrecks pose unique challenges for radiographic detection, particularly when hazardous materials are involved. crude oil and diesel fuel, stored in tanks, show moderate muon attenuation with relatively little scattering, making them relatively straightforward to detect. Explosives cause stronger attenuation combined with moderate scattering, complicating imaging. Steel-cased bombs, by contrast, create significant scattering and attenuation due to their high density. Voids or collapsed compartments appear as bright, low-density anomalies. Well-known examples of wrecks with hazardous cargo include the World War II ammunition ship S.S. Richard Montgomery, the crude oil tanker Montebello, and the Lusitania, long rumored to have carried munitions. In the Baltic region, many wrecks are suspected to contain either crude oil or chemical munitions, but their internal condition remains uncertain.
@@
Traditional Methods for Shipwreck Exploration
@@
As a result, traditional approaches can document external morphology and site context, but they remain limited when it comes to non-invasively mapping the internal structure, cargo, or hazardous materials of wrecks. This motivates the exploration of complementary approaches such as muon tomography, which can image density distributions through thick sediment and metallic hulls without physical disturbance.
Underwater Muon Imaging for Shipwreck Exploration
The method is particularly suited to scanning large dense objects. Muon attenuation signatures when passing through this objects can reveal crude oil and fuel tanks, explosives, or steel-cased munitions [17, 18, 19, 20]. Feasibility has already been demonstrated in Tokyo Bay, where detectors imaged through tens of metres of water and sediment [21, 22]. Similar deployments could be adapted for Baltic wrecks, generating either 2D transmission maps or full 3D tomographic reconstructions. Advances in compact scintillator arrays, readout electronics, and pressure housings make such systems increasingly feasible [23, 24], while simulations with Geant4 provide guidance for geometry optimization and exposure time [25].
Since muon flux decreases with depth, practical underwater muography is expected to be effective at depths up to approximately 100 m. Within this range, several examples of crude oil-carrying wrecks illustrate the potential applications:
Shipwrecks with crude oil: - The SS Montebello (sunk 1941 off California) lies in 270 m of water and still contains thousands of tonnes of crude crude oil, making it a well-studied case of environmental risk. Although deeper than typical muographic range, it demonstrates the scale of the problem. - The SS Richard Montgomery (1944), though located in shallow waters of the Thames Estuary, is loaded with munitions that pose an ongoing hazard and highlight the value of non-invasive imaging. - In the Baltic, numerous World War II-era tankers and freighters are suspected to retain significant crude oil or chemical munitions. Many lie in depths of 50–100 m, placing them within the feasible operating range of underwater muography.
Shipwrecks with munitions: SS Richard Montgomery (1944) – A U.S. Army cargo ship laden with munitions, sunk off the coast of England near Kent. The wreck is at approximately 30-70 meters, but parts or related sites could be near 100 meters. The German WWII ships (various) – Many sank with munitions, some around 100 meters depth in the North Sea or Baltic Sea.
Shipwrecks with archaeological artifacts, gold and treasures: Another potential application of underwater muography is the investigation of treasure-laden wrecks. Historical Spanish galleons, such as the Nuestra Señora de Atocha (sunk in 1622 off the Florida Keys), carried large quantities of gold and silver coins, jewelry, and other precious cargo. While the Atocha itself lies at a depth of about 40–50 meters, other galleons and treasure ships are reported at depths approaching 100 meters, where conventional diving surveys are challenging and sonar imaging provides limited information on buried or sediment-covered structures. In such cases, muon radiography could provide unique non-invasive insights by distinguishing high-density metallic inclusions (e.g., gold or silver coins) from surrounding sediments and hull remains.
@@ 2.1 Monte Carlo simulations
Figure 1: (a) Illustration of the tracking principle in muon telescope as simulated using GEANT4 code. (b) The schematic drawing illustrates muography imaging of an underwater shipwreck, simulated as a trapezoidal shape, with a muon detector positioned on the seabed and oriented toward the shipwreck.
@@
In GEANT4 [26], the detector is modeled as a set of three planes, each measuring $1\times 1\,\mathrm{m}^{2}$, with a typical vertical spacing of 0.5 m. This configuration functions as a muon telescope for tracking muons and reconstructing internal density via muon tomography. Each detector plane consists of two adjacent modules, totaling 64 plastic scintillator bars with triangular cross-sections and central holes for wavelength-shifting (WLS) fibers. The triangular geometry enables compact, crack-free arrays, while combining signals from adjacent bars improves spatial resolution through weighted position averaging. WLS fibers collect scintillation light and transmit it to silicon photomultipliers (SiPMs). Each plane provides one spatial coordinate (X or Y) of the muon hit. Two planes with orthogonal bars form an X-Y station, and the complete muon telescope consists of three such stations (Figure 1(a)). Muon trajectories are reconstructed by fitting straight lines through the recorded X-Y-Z hit positions across the three stations. The three-layer redundancy enhances track accuracy, reduces accidental triggers from SiPM noise, and improves cosmic-ray shower rejection. Future upgrades may include time-of-flight measurements to determine muon direction.
@@
Although this detector concept has been successfully applied in terrestrial settings, such as imaging pyramids and inspecting cargo containers, it is now being adapted for underwater deployment. In marine environments, detectors are positioned near or within wrecks to detect muons passing through structures, enabling noninvasive 3D imaging of internal components. To enhance detection efficiency, the detector is inclined at a 10-degree angle relative to the wreck.
In this study, we simulated a shipwreck as a trapezoidal object approximately 30 meters long, 15 meters high, and 15 meters wide, submerged at a depth of 50 meters to mimic the marine environment (see Figure 1(b)). The wreck was filled with either crude oil (density = 0.92 g/cm3) or hexogen (RDX) explosive (density = 1.82 g/cm3), depending on the scenario.
2.2 Modeling muon transport underwater
To analyze the contents of shipwrecks, such as determining whether they are filled with water, hazardous crude oil, or explosives, the muon flux transmission through the target can be measured. By examining the attenuation of muon flux as it passes through different materials, the muon transmission (MT) technique enables the creation of two- and three-dimensional average transmission and density maps of the studied object, providing valuable insights into internal relative densities. Achieving a well-defined muographic image of large targets, such as underwater shipwrecks, often requires data acquisition periods extending to weeks or even months. This extended duration is necessary to accumulate sufficient muon statistics across various viewing angles, ensuring adequate resolution and contrast in the resulting density maps.
The required exposure time depends on several factors, including the depth of the shipwreck, the size and density of the target, the detector’s acceptance and efficiency, and the desired spatial resolution.
The GEANT4 package was used to simulate muon propagation through water, from the sea surface down to the shipwreck location. Muons were generated at sea level with the CRY event generator [27], sampled over a $50\times 50$ m2 surface. Muons were propagated 35 meters underwater and detected by a virtual detector with dimensions of $20\times 20$ m2. To optimize computational resources, we adopted a strategy similar to that used for modeling underground muon fluxes, where the MUSIC code is employed to calculate muon transport through rock and the resulting distributions are then used by MUSUN to simulate muon flux in the underground laboratory and detector interactions. In our approach, GEANT4 was used to propagate muons through water, and the detector recorded only essential information: particle type (PDG code), kinetic energy, interaction position, and angular directions ($x$, $y$, $z$). These data provided tabulated underwater muon energy and angular distributions, which were then supplied to the PrimaryGeneratorAction.cc in GEANT4 to generate muons directly at a depth of 35 meters. The regenerated muons served as primary particles in different filling scenarios to model their interactions with a parameterized shipwreck geometry, enabling systematic studies of muon attenuation, scattering, and transmission imaging for internal structure reconstruction.

### v5

@@ Abstract
Muon imaging, a non-invasive technique that utilizes naturally occurring cosmic muons, has emerged as a promising tool for exploring underwater objects, including shipwrecks. This study investigates the potential of muon radiography to examine the contents of wrecked ships in the Baltic Sea and other marine environments. These wrecks often pose significant environmental risks due to hazardous contents such as explosives and crude oil, making their detection and monitoring critical for environmental and safety considerations. Accurate modeling and imaging of such wrecks are therefore essential for assessing potential dangers and mitigating environmental impacts.
To model the underwater muon flux in this study, an approach similar to that used in underground muon experiments was adopted. Cosmic muons were propagated through seawater using GEANT4, modeling their energy loss as they traverse water. The muons’ kinematics and energy depositions were recorded during this process. The resulting distributions were tabulated and used as input for the primary particle generation module. This method enabled the generation of a realistic underwater muon flux at the desired depth above the shipwreck, allowing us to simulate various shipwreck filling scenarios.
Assuming an exposure time of one week for a wreck located at 50 meters depth, our simulations demonstrate that muon imaging can sufficiently resolve density contrasts to distinguish between water, oil, and high-density materials. These results demonstrate the feasibility of muon radiography as a practical tool for underwater hazard assessment and shipwreck investigation.
@@ 1 Introduction
The deep sea holds countless sunken vessels, some of which may contain hazardous materials such as oil or explosives, making direct investigation risky for both safety and the preservation of the wreck. A promising, non-invasive solution comes from particle physics: muon imaging. Also known as muon tomography or radiography, this technique exploits naturally occurring cosmic-ray muons, highly penetrating particles that can reveal the internal structure of large, dense objects. Initially developed for applications such as scanning pyramids and monitoring nuclear reactors, muon imaging has since expanded to geophysics, archaeology, civil engineering and border security [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. Muon imaging has also been explored for underwater applications, with pilot studies demonstrating its considerable potential. Researchers at the University of Tokyo have successfully employed muon radiography to image underwater volcanic structures, showcasing its feasibility in marine environments. This technique involves deploying muon detectors beneath the water to visualize subsurface geological features. A notable example is the Tokyo Bay Seafloor Hyper-Kilometric Submarine Deep Detector (TS-HKMSDD), which is the world’s first under-seafloor muon detector array. Installed beneath the Tokyo Bay seafloor, this system collects muon data to monitor tidal variations and other subsurface phenomena[12]. Additionally, the University of Tokyo’s International Muography Research Organization has developed the wireless muometric navigation system (MuWNS), which leverages cosmic-ray muons to enable navigation in underground environments where GPS signals are unavailable[13]. These advancements highlight the promising potential of muon radiography as a non-invasive technique for underwater exploration and monitoring. In this context, its application to submerged environments offers exciting possibilities. By utilizing muon imaging to examine shipwrecks, researchers can analyze internal structures and contents without the need for excavation. This approach not only aids in the preservation of cultural heritage but also opens new avenues for uncovering historical and scientific insights beneath the waves.
@@
Underwater shipwrecks pose unique challenges for radiographic detection, particularly when hazardous materials are involved. Crude oil and diesel fuel, stored in tanks, show moderate muon attenuation with relatively little scattering, making them relatively straightforward to detect. Explosives cause stronger attenuation combined with moderate scattering, complicating imaging. Steel-cased bombs, by contrast, create significant scattering and attenuation due to their high density. Voids or collapsed compartments appear as bright, low-density anomalies. Well-known examples of wrecks with hazardous cargo include the ammunition ship S.S. Richard Montgomery, the crude oil tanker Montebello, and the Lusitania, long rumored to have carried munitions. In the Baltic region, many wrecks are suspected to contain either crude oil or chemical munitions, but their internal condition remains uncertain.
@@
Traditional methods for shipwreck exploration.
@@
As a result, traditional approaches can document external morphology and site context, but they remain limited when it comes to non-invasively mapping the internal structure, cargo, or hazardous materials of wrecks. This motivates the exploration of complementary approaches such as muon radiography, which can image density distributions through thick sediment and metallic hulls without physical disturbance.
Underwater muon imaging for shipwreck exploration.
The method is particularly suited to scanning large dense objects. Muon attenuation signatures when passing through this objects can reveal crude oil and fuel tanks, explosives, or steel-cased munitions [16, 17, 18]. Feasibility has already been demonstrated in Tokyo Bay, where detectors imaged through tens of meters of water and sediment [19]. Similar deployments could be adapted for Baltic wrecks, generating either 2D transmission maps or full 3D tomographic reconstructions. Advances in compact scintillator arrays, readout electronics, and pressure housings make such systems increasingly feasible [20], while simulations with GEANT4 provide guidance for geometry optimization and exposure time.
@@ 2.1 Monte Carlo simulations
In GEANT4 [21], the detector is modeled as a set of three planes, each measuring $1\times 1\,\mathrm{m}^{2}$, with a typical vertical spacing of 0.5 m. This configuration functions as a muon telescope for tracking muons and reconstructing internal density via muon radiography. Each detector plane consists of two adjacent modules, totaling 64 plastic scintillator bars with triangular cross-sections and central holes for wavelength-shifting (WLS) fibers. The triangular geometry enables compact, crack-free arrays, while combining signals from adjacent bars improves spatial resolution through weighted position averaging. WLS fibers collect scintillation light and transmit it to silicon photomultipliers (SiPMs). Each plane provides one spatial coordinate (x or y) of the muon hit. Two planes with orthogonal bars form an x-y station, and the complete muon telescope consists of three such stations (Figure 1). Muon trajectories are reconstructed by fitting straight lines through the recorded (x,y,z) hit positions across the three stations. The three-layer redundancy enhances track accuracy, reduces accidental triggers from SiPM noise and natural radioactivity.
@@
In marine environments, detectors are positioned near wrecks to detect muons passing through structures, enabling noninvasive 3D imaging of internal components. To enhance detection efficiency, the detector is inclined at a 10-degree angle relative to the wreck.
2.1.1 Modeling optical photons transport
We have developed a customizable simulation code for modeling optical photon transport in organic scintillators based on GEANT4. The code allows for the implementation of various geometries, materials, and reflective surfaces.
The scintillator bar is 1000 mm long, have a triangular cross-section (as shown in Figure 2(a)) with a base width of $33~\mathrm{mm}$, a height of $17~\mathrm{mm}$, and an extrusion hole of $2~\mathrm{mm}$ in diameter to accommodate the wavelength-shifting (WLS) fiber. The EJ-200 plastic scintillator [22] was selected as the base material. A scintillation yield of $10{,}000~\text{photons/MeV}$ was assumed to benchmark the optical model.
Extruded plastic scintillator bars are co-extruded with a white outer coating composed of 15% of titanium dioxide (TiO2) mixed with polystyrene. This coating functions as a diffusive internal reflector for scintillation light. By reflecting and directing scintillation photons toward the embedded wavelength-shifting fibers within the scintillator core, the TiO2 coating enhances light collection efficiency. The co-extrusion process for applying this reflective layer is more cost-effective than alternative methods, such as wrapping the scintillator with reflective materials, and it provides a more uniform, air-free coating. The light attenuation length of scintillator bar was L= 30 cm [23].
The scintillator strips are read out by 1.5 mm diameter Kuraray multiclad WLS fibers, 11-Y (S35 type) [24]. The SiPMs detect scintillation light from both ends of the WLS fibers. To reduce light losses, an optical coupler was applied in a 2.0 mm diameter hole, which formed the gap between the fiber and the scintillator bars. The scintillator bars have an index of refraction $n_{\text{sc}}=1.59$, while the WLS fibers have indices of refraction $n_{\text{WLS}}=1.59$, 1.49, and 1.42 for the fiber core, inner cladding, and outer cladding, respectively. The silicone optical grease used as the coupler in the gap had a refractive index of $n_{\text{gap}}=1.50$, which optimizes light transmission and reflection properties. Hamamatsu S13360-2050VE devices [25] with a 2 × 2 mm2 photosensitive area is used to detect optical photons. The SiPMs are modeled as a Si material. To improve light detection, SiPMs are coupled with WLS fiber with an Silicone Optical Grease BC-631 [26] with a refractive index of 1.465. The light detection efficiency is determined by weighting the SiPM PDE with the Kuraray WLS fiber emission spectra, which gives a value of $\approx$40%.
In this work, the glisur model is employed the, which accounts for surface roughness (polishing degree) through a roughness parameter in the range 0.5–0.9. The surfaces are modeled as dielectric_dielectric, co-extruded coating was treated as dielectric_metal surface, where photons are reflected with a probability corresponding to the set reflectivity, non-reflected photons are absorbed.
Figure 1: (a) GEANT4 simulation of optical photon propagation in a single layer of stacked triangular bars is shown. (b) Two orthogonally oriented layers form a $1\times 1~\mathrm{m}^{2}$ superlayer used for determining the xxyy coordinates of the muon hit position.
2.1.2 Position resolution estimation
To evaluate the hit position, muon beams were simulated with a 2 mm step from position $X_{1}$ to $X_{2}$. For each beam position, 1000 muons were generated. The difference in the number of detected optical photons between the upper and lower triangular scintillation bars provides sensitivity to the muon impact position. When two scintillation bars record signals, the muon impact position along the detector axis can be estimated from the ratio of detected photons, $N_{2}/N_{1}$, in the upper and lower bars. A geometric correction factor accounting for the track inclination angle $\vartheta$ is applied [27]:
$x=x_{1}+\frac{a\,N_{2}}{N_{1}+N_{2}}+h\frac{N_{2}}{N_{1}+N_{2}}\tan\vartheta$ (2.1)
where $a$ is the fiber spacing ($a$ = $X_{1}$-$X_{2}$), $h$ the rod height, and $N_{1}$, $N_{2}$ the detected photons in upper and lower super layers. Figure 2(b) shows the distribution of reconstructed hit positions for two muon beams separated by 2 mm. The two peaks correspond to beam positions at 3 and 5 mm (assuming X1 = 0), demonstrating that the light-yield asymmetry between the upper and lower triangular scintillation bars provides a sensitive measure of the muon impact point.
Figure 2: (a) Cross-sectional view of the triangular scintillator bars illustrating the geometry used for hit position reconstruction. The red lines, separated by 2 mm, represent two incident muon trajectories simulated to evaluate the detector’s spatial resolution. (b) Distribution of reconstructed hit positions using asymmetry between upper and lower triangular scintillation bars for two simulated muon beams separated by 2 mm.
The number of optical photons detected in each layer was used to calculate the hit position according to Equation 2.1. The position resolution, $\sigma_{x}$, can be estimated from the overlap region of the two distributions by fitting them with Gaussian functions and evaluating their standard deviations. This method enables the use of data from a pair of orthogonal layers (a superlayer) to determine the local hit position in both x and y coordinates.
The ability of the reconstruction algorithm to distinguish closely spaced muon trajectories was evaluated using a Receiver Operating Characteristic (ROC) analysis. The ROC curve was generated from the distribution of reconstructed hit positions, obtained from the asymmetry between the upper and lower triangular scintillator bars, for two simulated muon beams separated by 2 mm. The resulting area under the curve (AUC) of 0.993, as shown in Figure 3, indicates excellent discrimination capability, confirming that the detector geometry and reconstruction method provide sufficient spatial resolution to reliably separate muon tracks with millimeter-scale spacing.
Figure 3: The ROC curve for the distribution of reconstructed hit positions, obtained from the asymmetry between the upper and lower triangular scintillator bars for two muon trajectories separated by 2 mm.
2.1.3 Muon track reconstruction
The muon trajectory is discretized as it passes through scintillation layers, with the impact point called an hit position. The telescope has three detection planes. Each plane comprises two orthogonal superlayers; each superlayer consists of triangular scintillator bars (strips) (as shown in Figure 4). A muon crossing a plane typically deposits energy in several neighbouring triangular strips of a superlayer. For the geometry described, a muon typically produces up to four fired strips per plane (two in each orthogonal layer), so that across the three planes 12 strip register signals. In realistic simulations, optical photon tracking for such a detector geometry would require simulating tens of thousands of photons per muon, a simplified approach is adopted. Each triangular scintillator bar is defined as a sensitive detector, and the energy deposition and hit position coordinates are recorded for every muon that passes through it. Instead of using the number of optical photons detected, the energy deposition in each layer was used to calculate the hit position according to Equation 2.1. This method allows data from a pair of orthogonal layers (a superlayer) to determine the local hit position in both $x$ and $y$ coordinates with high precision.
In this approach, the degradation of position resolution due to Landau fluctuations in the $dE/dx$ energy loss, Poisson fluctuations in light collection (and photoelectron generation), and spatially dependent non-uniformities in light collection is not explicitly included. To account for these effects, the hit coordinates $(x,y)$ in each detection plane are smeared using a Gaussian distribution with a standard deviation chosen to reproduce a realistic spatial resolution of 2 mm. The final muon trajectory is then reconstructed by fitting these smeared hit positions across the three detector modules (superlayers).
Figure 4: Illustration of the muon tracking principle in a telescope, from GEANT4 simulation. Muons with positive charge are shown in red and muons with negative charge in blue. A muon track is reconstructed when a spatial coincidence of hits is observed in all three tracking planes. Scintillation bars activated by the passing muon are highlighted in cyan.
2.1.4 Muons event generator for modeling muon transport underwater
To optimize computational resources, we adopted a strategy similar to that used for modeling underground muon fluxes, where the MUSIC (MUon SImulation Code) code is employed to calculate muon transport through rock and the resulting distributions are then used by MUSUN (MUon Simulations UNderground) to simulate muon flux in the underground laboratory and detector interactions [28]. MUSIC and MUSUN are essential tools for simulating cosmic-ray muons in underground research. MUSIC acts as a muon propagator, modeling energy loss and scattering as muons traverse rock or water overburden. It provides detailed energy and angular distributions of muons at various depths. MUSUN uses MUSIC’s output to generate realistic muon populations, defining their position, energy, and direction, as they would appear in or around an underground detector. Together, MUSIC and MUSUN enable accurate predictions of underground muon fluxes, crucial for experiment design and data interpretation in particle and astroparticle physics. By integrating accurate muon transport modeling with realistic muon generation, these tools enable the mitigation of muon-induced backgrounds, improving the sensitivity of dark matter, neutrinoless double beta decay, and other rare event searches. Alternatively, the dedicated Monte Carlo codes (MUon inTensity codE) and PROPOSAL [29], MUPAGE (MUon GEnerator from PArametric formulas) [30] simulate muon transport and generate muons at underground sites.
In our approach, the GEANT4 package was used instead of MUSIC to simulate muon propagation through water, from the sea surface down to the shipwreck location. Muons were generated at sea level with the CRY event generator [31], sampled over a 50 $\times$  50 m2 surface. Muons were propagated through 35 meters of water and interacted with a sensitive surface measuring 20 $\times$ 20 m2. Out of 109 generated muons, approximately 1.17 $\cdot$ 107 reached this surface.
For these muons essential information was recorded: particle type (PDG code), kinetic energy, interaction position, and muon momentum direction cosines ($u_{x}$, $u_{y}$, $u_{z}$). Similar to the MUSUN code, the tabulated underwater muon energy and angular distributions were implemented in the PrimaryGeneratorAction class in GEANT4, allowing muons to be generated directly at a depth of 35 meters. In GEANT4, the PrimaryGeneratorAction class (or a class derived from G4VUserPrimaryGeneratorAction) is a user-defined class that defines how primary particles are generated for a GEANT4 simulation. It is one of the three mandatory classes required for a GEANT4 application and is responsible for creating primary events by setting up particles, their initial positions, energies, and directions. For simple distributions, GEANT4 provides the capability to define particle generators based on mathematical, analytical formulas. For more complex or non-standard distributions, such as those found in real experimental data, it is possible to load data from external files. This involves reading distribution information from data tables, which can then be used to generate primary particles with the desired physical properties. For example, in the GEANT4 generic examples, the FCALPrimaryGeneratorAction class implements a particle generator that reads particle data from external files. In our approach, the external file contains data on muon charge, position, energy, and angular distributions at a depth of 35 m. These data are used to generate muons at this depth, serving as primary particles in various filling scenarios to model their interactions with a parameterized shipwreck geometry.